If you can earn 8 percent annual return on your investments, how long does it take to quadruple (4 times) your investments?
Answer #1

Solution :

Let's suppose our investent = X

Final value = 4X ( 4 times the investment )

Interest rate = 8% per annum

Using Future value formula

Final value = initial investment * ( 1+interest rate) ^ year

4X = X * ( 1+8%) ^ year

1.08 ^ year = 4

Taking log both side

log ( 1.08^ year ) = log 4

Year * log 1.08 = log 4

Year = log 4 / log 1.08 = 0.602/0.0334 = 18

So correct answer is 18 years